# The lattice structure of negative Sobolev and extrapolation spaces

### v1

@@ 4. Extrapolation spaces of positive semigroups
Let $X$ be a Banach lattice and let $(T(t))_{t\geq 0}$ be a $C_{0}$-semigroup on $X$ that is positive, which means that each operator $T(t)$ is positive. Let $A$ denote the generator of $(T(t))_{t\geq 0}$; which is a closed and densely defined operator. For the general theory of $C_{0}$-semigroups, we refer to [6] and for positive $C_{0}$-semigroups specifically to [12, 5].
Fix a real number $\lambda$ that is larger than the real part of every spectral value of $A$. The extrapolation space $X_{-1}$ is defined as follows: for every $x\in X$ one sets $\left\lVert x\right\rVert_{-1}:=\left\lVert(\lambda-A)^{-1}x\right\rVert_{X}$ and then defines $X_{-1}$ to be the completion of $X$ with respect to this norm. Note that $X_{-1}$ does not depend on the choice of $\lambda$ since all norms that one obtains in this way are equivalent. One can extend the semigroup generator $A$ to an operator $A_{-1}:X\subseteq X_{-1}\to X_{-1}$. The resolvent can also be extended to an operator from $X_{-1}$ to $X$ that then intertwines $A$ and $A_{-1}$.
The space $X_{-1}$ plays, for instance, a role in perturbation and linear systems theory. In order to study positive perturbations the authors of [4] introduced a cone $X_{-1,+}$ in $X_{-1}$ which they defined to be the closure of $X_{+}$ within $X_{-1}$. This is indeed a cone and satisfies $X_{-1,+}\cap X=X_{+}$ [4, Proposition 2.3]. The extended resolvent operators $(\mu-A_{-1})^{-1}$ are positive from $X_{-1}$ to $X$ for each $\mu$ that is larger than the real part of all spectral values of $A$ [4, Remark 2.2]. Various order properties of the cone $X_{-1,+}$, even in the case that $X$ is an ordered Banach space were recently proved in [3, Section 2.2].
@@
Let $(T(t))_{t\geq 0}$ be a positive $C_{0}$-semigroup on a Banach lattice $X$ with order continuous norm. Then $\operatorname{span}(X_{-1,+})$ is – with respect to a norm equivalent to $\left\lVert\mathord{\,\cdot\,}\right\rVert_{\operatorname{span}X_{-1,+}}$ – a Banach lattice with order continuous norm. If $X$ is a KB-space, then so is $\operatorname{span}(X_{-1,+})$.
@@
In some situations one can concretely determine the space $\operatorname{span}X_{-1,+}$:
@@
Let $\Omega\subseteq\mathbb{R}^{d}$ be a bounded domain with Lipschitz boundary, let $p,q\in(1,\infty)$ be Hölder conjugates, and let $\Delta$ denote the Neumann Laplace operator on $X=L^{p}(\Omega)$. If $d<2q$, then $\operatorname{span}X_{-1,+}$ is precisely the space of finite Borel measures on $\overline{\Omega}$; see [3, Example 2.13(b)]. Note that, since positive operators on ordered Banach spaces with generating cone are automatically continuous [1, Theorem 2.32], all complete norms that turn $\operatorname{span}X_{-1,+}$ into an ordered Banach space, are equivalent. Hence, the norm $\left\lVert\mathord{\,\cdot\,}\right\rVert_{\operatorname{span}X_{-1,+}}$ is equivalent to the total variation norm in this example and consequently, the space $\operatorname{span}X_{-1,+}$ is not reflexive despite $X$ and $Z$ both being reflexive.
Since the assumptions of Corollary 2.4 are also satisfied in this example (with $Z:=X_{-1}$) this also shows that the conclusion of Corollary 2.4 cannot be strengthened to give that $\operatorname{span}Z_{+}$ is reflexive.
@@ 5. Outlook
Assume that we are in the situation of Theorem 2.5, but that the Banach lattice $X$ is not required to have order continuous norm. We do not know whether $\operatorname{span}Z_{+}$ still needs to be a Banach lattice in this case (the arguments in the proof of Theorem 2.5 show that it suffices to prove that $\operatorname{span}Z_{+}$ is a vector lattice). This is the first of the following two open problems.
@@
[…]
    Let $X$ be a Banach lattice and let $Z$ be an ordered Banach space. Assume that there exists a positive bijection $S\in\mathcal{L}(Z,X)$, a positive operator $J\in\mathcal{L}(X,Z)$ and a sequence of positive operators $(R_{n})\subseteq\mathcal{L}(Z,X)$ such that $(JR_{n})\subseteq\mathcal{L}(Z)$ converges to the $\operatorname{id}_{Z}$ with respect to the strong operator topology.
    Does it follow that $\operatorname{span}Z_{+}$ is a vector lattice (and thus a Banach lattice with respect to a norm equivalent to $\left\lVert\mathord{\,\cdot\,}\right\rVert_{\operatorname{span}Z_{+}})$?
[…]
    If the answer to problem (a) is negative, does it become positive if the mapping $J$ is compact?
As explained before the problem statement, Theorem 2.5 says that the answer to problem (a) is positive if $X$ has order continuous norm. Two concrete examples in which $X$ does not have order continuous norm are given in [4, Examples 5.1 and 5.3]. In both examples, $X$ is a space of continuous functions on an interval (so it does not have order continuous norm) and $Z:=X_{-1}$ is the extrapolation space of a positive $C_{0}$-semigroup on $X$. However, in both examples the positive cone is computed explicitly in [4, formulas (5.4) and (5.6)] and those results show that the span of the cone is indeed a lattice. In both examples the canonical embedding from $X$ into $X_{-1}$ is compact (since the generator of the semigroup has compact resolvent in both cases). This might be an indication that the answer to problem (b) above is positive.
@@ Acknowledgements
The first author was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – 523942381. The second author is indebted to the Department of Applied Mathematics of the University of Twente for a very pleasant stay during which some parts of the work on this article were done. The authors are grateful to Lassi Paunonen for useful comments on the manuscript.

### v3

@@ 4. Extrapolation spaces of positive semigroups
Let $X$ be a Banach lattice and let $(T(t))_{t\geq 0}$ be a $C_{0}$-semigroup on $X$ that is positive, which means that each operator $T(t)$ is positive. Let $A$ denote the generator of $(T(t))_{t\geq 0}$; this is a closed and densely defined operator. For the general theory of $C_{0}$-semigroups, we refer to [7] and for positive $C_{0}$-semigroups specifically to [15, 6].
Fix a real number $\lambda$ that is larger than the real part of every spectral value of $A$. The extrapolation space $X_{-1}$ is defined as follows: for every $x\in X$ one sets $\left\lVert x\right\rVert_{-1}:=\left\lVert(\lambda-A)^{-1}x\right\rVert_{X}$ and then defines $X_{-1}$ to be the completion of $X$ with respect to this norm. Note that $X_{-1}$ does not depend on the choice of $\lambda$ since all norms that one obtains in this way are equivalent. One can extend the semigroup generator $A$ to a unique operator $A_{-1}\in\mathcal{L}(X,X_{-1})$. The resolvent can also be extended to an operator from $X_{-1}$ to $X$ that then intertwines $A$ and $A_{-1}$.
The space $X_{-1}$ plays, for instance, an important role in perturbation and linear systems theory. In order to study positive perturbations and positive inputs in systems theory, a natural cone is needed in the extrapolation space $X_{-1}$. Following [5, 18] we define the cone $X_{-1,+}$ in $X_{-1}$ as
$X_{-1,+}:=\overline{X_{+}}^{X_{-1}},$ (4.1)
i.e., as the closure of $X_{+}$ within $X_{-1}$. This is indeed a cone and it satisfies $X_{-1,+}\cap X=X_{+}$ [5, Proposition 2.3]. The extended resolvent operators $(\mu-A_{-1})^{-1}$ are positive from $X_{-1}$ to $X$ for each $\mu$ that is larger than the real part of all spectral values of $A$ [5, Remark 2.2]. Various order properties of the cone $X_{-1,+}$, even in the case that $X$ is an ordered Banach space, were recently proved in [3, Section 2.2].
@@
Let $(T(t))_{t\geq 0}$ be a positive $C_{0}$-semigroup on a Banach lattice $X$ with order continuous norm. Then $\operatorname{span}(X_{-1,+})$ is a Banach lattice with order continuous norm with respect to a norm equivalent to $\left\lVert\mathord{\,\cdot\,}\right\rVert_{\operatorname{span}(X_{-1,+})}$. Moreover, the canonical embedding of $X$ into $\operatorname{span}(X_{-1,+})$ is a lattice homomorphism and $X$ is a lattice ideal in $\operatorname{span}(X_{-1,+})$. If $X$ is a KB-space, then so is $\operatorname{span}(X_{-1,+})$.
@@
In some situations one can concretely determine the space $\operatorname{span}(X_{-1,+})$:
@@
Let $\Omega\subseteq\mathbb{R}^{d}$ be a bounded domain with Lipschitz boundary, let $p,q\in(1,\infty)$ be Hölder conjugates, and let $\Delta$ denote the Neumann Laplace operator on $X=L^{p}(\Omega)$. If $d<2q$, then $\operatorname{span}(X_{-1,+})$ is precisely the space of finite Borel measures on $\overline{\Omega}$; see [3, Example 2.13(b)]. Note that, since positive operators on ordered Banach spaces with generating cone are automatically continuous [1, Theorem 2.32], all complete norms that turn $\operatorname{span}(X_{-1,+})$ into an ordered Banach space, are equivalent. Hence, the norm $\left\lVert\mathord{\,\cdot\,}\right\rVert_{\operatorname{span}(X_{-1,+})}$ is equivalent to the total variation norm in this example and consequently, the space $\operatorname{span}(X_{-1,+})$ is not reflexive despite $X$ and $X_{-1}$ both being reflexive.
Since the assumptions of Corollary 2.6 are also satisfied in this example (with $Z:=X_{-1}$) this also shows that the conclusion of Corollary 2.6 cannot be improved to give that $\operatorname{span}(Z_{+})$ is reflexive.
Let us conclude this section with another example of an extrapolation space, which is simpler but nevertheless illuminating.
Example 4.3.
Let $p\in[1,\infty)$, let $(\Omega,\mu)$ be a measure space, and endow $L^{p}(\Omega,\mu)$ with the pointwise almost everywhere order. Let $m:\Omega\to[0,\infty)$ be measurable. The operator $A:f\mapsto-mf$ on $L^{p}(\Omega,\mu)$ with domain
$\operatorname{dom}\left(A\right):=\{f\in L^{p}(\Omega,\mu):\,mf\in L^{p}(\Omega,\mu)\}$
generates a positive $C_{0}$-semigroup $(T(t))_{t\geq 0}$ given by
$T(t)f=e^{-tm}f$
for each $f\in L^{p}(\Omega,\mu)$ and each $t\in[0,\infty)$. It is easy to check that the extrapolation space $\big{(}L^{p}(\Omega,\mu)\big{)}_{-1}$ equals $L^{p}(\Omega,\nu)$ with $\mathrm{d}\nu=\frac{1}{(m+1)^{p}}\mathrm{d}\mu$, and that the cone (4.1) in the extrapolation space consists simply of the functions in $L^{p}(\Omega,\nu)$ that are $\geq 0$ $\nu$-almost everywhere.
In contrast to Example 4.2, the cone $X_{-1,+}$ in the extrapolation space $X_{-1}:=\big{(}L^{p}(\Omega,\mu)\big{)}_{-1}$ in Example 4.3 is generating, so $\operatorname{span}(X_{-1,+})=X_{-1}$.
@@ 5. Outlook
Assume that we are in the situation of Theorem 2.8, but that the Banach lattice $X$ is not required to have order continuous norm. We do not know whether $\operatorname{span}(Z_{+})$ still needs to be a Banach lattice in this case (the arguments in the proof of Theorem 2.8 show that it suffices to prove that $\operatorname{span}(Z_{+})$ is a vector lattice). This is the first of the following two open problems.
@@
(a) Let $X$ be a Banach lattice and let $Z$ be an ordered Banach space. Assume that there exists a positive bijection $S\in\mathcal{L}(Z,X)$, a positive operator $J\in\mathcal{L}(X,Z)$ and a sequence of positive operators $(R_{n})\subseteq\mathcal{L}(Z,X)$ such that $(JR_{n})\subseteq\mathcal{L}(Z)$ converges to the $\operatorname{id}_{Z}$ with respect to the strong operator topology. (If it is helpful, one might also want to assume that, in addition, $R_{n}J\to\operatorname{id}_{X}$ strongly.)
Does it follow that $\operatorname{span}(Z_{+})$ is a vector lattice (and thus a Banach lattice with respect to a norm equivalent to $\left\lVert\mathord{\,\cdot\,}\right\rVert_{\operatorname{span}(Z_{+})})$?
(b) If the answer to problem (a) is negative, does it become positive if the mapping $J$ is compact?
As explained before the problem statement, Theorem 2.8 says that the answer to problem (a) is positive if $X$ has order continuous norm. Two concrete examples in which $X$ does not have order continuous norm are given in [5, Examples 5.1 and 5.3]. In both examples, $X$ is a space of continuous functions on an interval (so it does not have order continuous norm) and $Z:=X_{-1}$ is the extrapolation space of a positive $C_{0}$-semigroup on $X$. However, in both examples the positive cone is computed explicitly in [5, formulas (5.4) and (5.6)] and those results show that the span of the cone is indeed a lattice. In both examples the canonical embedding from $X$ into $X_{-1}$ is compact (since the generator of the semigroup has compact resolvent in both cases). This might be an indication that the answer to problem (b) above is positive.
@@ Acknowledgements
The first author was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – 523942381. The second author is indebted to the Department of Applied Mathematics of the University of Twente for a very pleasant stay during which some parts of the work on this article were done. The authors are grateful to Lassi Paunonen for useful comments on the manuscript. The authors are also indebted to Jens Wintermayr for asking whether the reflexivity assumption in Corollary 2.6 can be weakened, a question that inspired Theorem 2.8.